# FALCON: Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN

### v2

@@ Abstract
Recently, large language models (LLMs) have achieved significant progress in automated code generation. Despite their strong instruction-following capabilities, these models frequently struggled to align with user intent in the coding scenario. In particular, they were hampered by datasets that lacked diversity and failed to address specialized tasks or edge cases. Furthermore, challenges in supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) led to failures in generating precise, human-intent-aligned code. To tackle these challenges and improve the code generation performance for automated programming systems, we propose Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN (i.e., FALCON). FALCON is structured into two hierarchical levels, from the global level, long-term memory improves code quality by retaining and applying learned knowledge, while from the local level, short-term memory allows for the incorporation of immediate feedback from compilers and AI systems. Additionally, we introduce meta-reinforcement learning with feedback rewards to solve the global-local bi-level optimization problem and enhance the model’s adaptability across diverse code generation tasks. Extensive experiments are conducted and it is found that our technique achieves state-of-the-art performance, leading other reinforcement learning methods by more than 4.5 percentage points on the MBPP benchmark and 6.1 percentage points on the Humaneval benchmark. The open-sourced code is publicly available at https://anonymous.4open.science/r/FALCON-BFE0/README.md.
1 Introduction
The development of Large Language Models (LLMs) has significantly advanced automated code generation (Zan et al., 2023). Models like CodeLLaMA (Roziere et al., 2023) and DeepSeek-Coder (Guo et al., 2024), tailored for code-centric tasks, have demonstrated outstanding performance across programming challenges. While LLMs excel in instruction-following through tuning (Jiang et al., 2024), they often misalign with user intent, making feedback-based adjustments critical. For example, InstructGPT (Ouyang et al., 2022) leverages reinforcement learning with human feedback (RLHF), and CodeRL (Le et al., 2022) uses compilation feedback to refine model performance. Similarly, CompCoder (Wang et al., 2022) enhances code compilability with compiler feedback, and RLTF (Liu et al., 2023a) offers fine-grained feedback on compiler errors. However, current RL frameworks generate compilation errors and overlook non-differentiable features (e.g. coding style) that affect the performance significantly (Jiang et al., 2024). To address these challenges, we propose a reinforcement learning system combining long-term and short-term memory feedback. From the global level, long-term memory tracks trends over time for higher-quality code retrieval, while from the local level, short-term memory captures recent errors and immediate feedback. The main contributions of this paper are as follows:
@@ 1 Introduction
  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to efficiently generalize across diverse programming tasks, adapt quickly to new environments, and handle a wide range of coding challenges with fewer training iterations.
2 Related Works
2.1 Pre-trained Models for Code Generation
@@
In recent years, pre-trained language models have made significant progress in the field of code generation. Trained on large-scale code corpora, these models have demonstrated powerful code generation capabilities. For example, CodeBERT Feng et al. (2020), a model based on an encoder-only architecture, has shown impressive results. With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling Vaswani et al. (2017). Several models, such as CodeGPT Lu et al. (2021), CodeGeeX Zheng et al. (2024), and DeepSeek-Coder Guo et al. (2024), use Causal Language Modelling (CLM) pretraining, while others like CodeT5 Wang et al. (2021) and AlphaCode Li et al. (2022) utilize an encoder-decoder architecture. Additionally, models like CodeFusion Singh et al. (2023) leverage diffusion-based techniques for code generation. These pre-trained models exhibit great potential in code generation tasks, achieving notable improvements in accuracy through various architectures and training strategies. However, they still face challenges in ensuring the syntactical and functional correctness of the generated code.
2.2 Reinforcement Learning on code
@@
Reinforcement learning (RL) is a method that learns optimal strategies through reward signals from interacting with the environment Fujimoto et al. (2019). It excels in sequence generation tasks, such as enhancing performance in translation and summarization models by improving BLEU and ROUGE scores Ahn et al. (2019). Unlike traditional natural language processing tasks, code generation is more complex. Beyond syntactical correctness, the generated code must be functionally accurate, meaning it should compile and perform as expected in various scenarios. Passing unit tests is important for verifying correctness, safety, and precision, but not sufficient unless the tests are comprehensive and well-designed. Unit tests must cover diverse use cases and edge cases to ensure the code adheres to standards and meets requirements. For instance, the RL-based code generation fine-tuning framework, CodeRL Le et al. (2022), guides fine-tuning by integrating unit test signals with reinforcement learning. PPOCoder Shojaee et al. (2023) improves on CodeRL by employing Proximal Policy Optimization to refine the approach, while RLTF Liu et al. (2023a) incorporates specific error information from the code and multi-granularity feedback alongside an online framework for model training. StepCoder Dou et al. (2024) enhances the code generation process through compiler feedback and segmental optimization. However, current RL-based methods still face some limitations. Primarily, these methods lack detailed feedback on specific types and distributions of programming errors, which is crucial for identifying patterns, understanding root causes, and creating targeted interventions. This makes addressing recurring issues difficult. Additionally, the corrective mechanisms are not robust enough, often failing to provide timely, specific guidance, hindering effective learning. Additionally, the diversity of input-output examples in existing benchmarks is limited, restricting the model’s ability to adapt and generalize to different or unseen problems Liu et al. (2023a).
3 Problem Setting
@@
We aim to enhance the automated code generation capabilities of Large Language Models (LLMs) by addressing key challenges in accuracy, diversity, error correction, and code quality. Formally, given a high-level specification $D$, the task is to generate a sequence of code tokens $W=\{w_{1},w_{2},\ldots,w_{T}\}$ that maximizes the conditional probability $P(W|D,\theta)$, where $\theta$ represents the model parameters. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\log P(W|D,\theta)\right]$. Challenges. Effective code generation is impeded by several factors. Accuracy requires minimizing syntactical or logical errors in $W$ to ensure correct functionality. Diversity of input-output examples $\mathcal{X}$ is often limited, restricting the model’s ability to generalize across varied programming tasks. Efficient error correction mechanisms are necessary to identify and rectify errors in $W$, ensuring robust performance. Maintaining high code quality, which encompasses adherence to coding standards, style guidelines, and managing code complexity $C(W)$, remains a persistent challenge. Additionally, adaptability refers to the model’s capacity to adapt to new tasks and incorporate feedback for continuous improvement, which is constrained without robust memory mechanisms. FALCON Framework. To address these challenges, we propose the FALCON framework. The framework leverages both a long-term memory buffer $\mathcal{M}_{\text{long}}=\{(D_{i},W_{i},T_{i},F_{i})\}_{i=1}^{N}$ and a short-term memory buffer $\mathcal{M}_{\text{short}}=\{(D_{j},W_{j},T_{j},F_{j})\}_{j=1}^{M}$ to utilize diverse feedback, enabling fine-tuning of $\theta$. Formally, FALCON seeks to optimize $\theta$ by maximizing a composite reward function $R(W,F)=\alpha T(W)+\beta S(W)+\gamma C(W)+\delta E(W)$, where $T(W)$, $S(W)$, $C(W)$, and $E(W)$ represent unit test, code style, complexity, and error feedback, respectively. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\mathbb{E}_{W\sim P(W|D,\theta)}\left[R(W,F)\right]\right]$. Assumptions. The effectiveness of FALCON is based on the following assumptions: 1. Exchangeability of the dataset $\mathcal{D}=\{D_{i}\}_{i=1}^{N}$, implying that the order of tasks does not influence model performance; 2. Independence of feedback signals $F$ given the generated code $W$; 3. Sufficient memory capacity in both $\mathcal{M}_{\text{long}}$ and $\mathcal{M}_{\text{short}}$ to store relevant interactions without significant data loss; and 4. Feedback efficacy, ensuring that integrated feedback mechanisms provide meaningful and actionable information to guide the optimization of $\theta$. Optimization Objective. Given a set of tasks $\mathcal{T}=\{D_{i}\}_{i=1}^{N}$, the goal is to learn optimal parameters $\theta^{*}$ that maximize the expected composite reward across all tasks, formulated as $\theta^{*}=\arg\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}R(W_{i},F_{i})$, where $W_{i}\sim P(W|D_{i},\theta)$ is the generated code for task $D_{i}$, and $F_{i}$ represents the aggregated feedback. The formulation ensures continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
4 Methodology
@@
4.1 Task Definition: Code Generation by LLMs
@@
4.2 MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
We propose the MRLF algorithm 1, which involves random task sampling from the task distribution during training. Previous works have indicated that different data sampling strategies have varying impacts on model information extraction Zhang et al. (2024). Consequently, we implement both long and short memory sequences. The long memory strategy stores the solutions generated for each problem and the compiler feedback results, whereas the short memory sequence selects the latest samples and unit test feedback from each current iteration. To address repetitive runtime and compilation errors, the long memory strategy categorizes and stores various errors by their types and records the corresponding error lines, enabling fine-grained reward allocation.
@@ 4.2 MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
1:Task distribution $\mathcal{T}$
2:Updated model parameters $\theta$
3:Initialize $LMB$, $SMB$, $\theta$
4:Populate $LMB$, $SMB$ via few-shot demonstrations
5:repeat
6:     Sample batch $\{T_{i}\}$ from $\mathcal{T}$
7:     for each task $T_{i}$ do
8:         Code Generation: Generate code $\hat{W}$ and test results; record in $SMB$
9:         Adaptation: Set $\theta_{i}=\theta$; sample mini-batch from $LMB$, $SMB$
@@
11:         Update $\theta_{i}$: $\theta_{i}\leftarrow\theta_{i}-\alpha\nabla_{\theta_{i}}L_{\text{inner}}(\theta_{i})$
12:         Evaluation: Assess $\theta_{i}$ on $T_{i}$; document outcomes
13:     end for
14:     Meta-Update: $L_{\text{meta}}=\sum_{i=1}^{N}L_{\text{inner}}(\theta_{i})$; refine $\theta$: $\theta\leftarrow\theta-\beta\nabla_{\theta}L_{\text{meta}}$
15:until Convergence
16:return $\theta$
@@
As depicted in Figure 2, we employ a meta-reinforcement learning framework to optimize code generation by integrating both long- and short-term memories for enhanced adaptability. From a global perspective, Long-Term Memory $D_{\text{long}}$ stores historical tasks, generated codes, and feedback to provide valuable context. From a local perspective, Short-Term Memory $D_{\text{short}}$ focuses on recent feedback to enable real-time adjustments. This approach leverages the MAML framework (Finn et al., 2017) for efficient task adaptation with minimal updates.
@@
4.3 Long-Term Memory Feedback
Retrieving information from long-term memory helps improve code quality. We use the FAISS framework (Douze et al., 2024) to retrieve relevant historical code, feedback, and evaluation scores. Task descriptions and feedback are transformed into embedding vectors and then indexed. During code generation, a query vector from the current task retrieves the top-k most similar historical data to guide the process and avoid past errors. The prompt template is provided in the appendix. Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score. We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\bm{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS. During the code generation phase, the current task description $\bm{t}_{\text{current}}$ and feedback $\bm{f}$ are transformed into a query vector $\bm{q}=\phi(\bm{t}_{\text{current}})$. We compute the similarity between the query vector $\bm{q}$ and the historical vectors $\bm{v}_{i}$ using cosine similarity $\text{cos}(\bm{q},\bm{v}_{i})$, and retrieve the top-$k$ most similar historical tasks. The retrieval process can be represented as:
$\{(t_{i_{1}},f_{i_{1}},e_{i_{1}}),\dots,(t_{i_{k}},f_{i_{k}},e_{i_{k}})\}=\text{Top-}k\left(\bm{v}_{i}\mid i=1,2,\dots,n\right)$ (7)

### v3

@@ Abstract
Recently, large language models (LLMs) have achieved significant progress in automated code generation. Despite their strong instruction-following capabilities, these models frequently struggled to align with user intent in the coding scenario. In particular, they were hampered by datasets that lacked diversity and failed to address specialized tasks or edge cases. Furthermore, challenges in supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) led to failures in generating precise, human-intent-aligned code. To tackle these challenges and improve the code generation performance for automated programming systems, we propose Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN (i.e., FALCON). FALCON is structured into two hierarchical levels, from the global level, long-term memory improves code quality by retaining and applying learned knowledge, while from the local level, short-term memory allows for the incorporation of immediate feedback from compilers and AI systems. Additionally, we introduce meta-reinforcement learning with feedback rewards to solve the global-local bi-level optimization problem and enhance the model’s adaptability across diverse code generation tasks. Extensive experiments are conducted and it is found that our technique achieves state-of-the-art performance, leading other reinforcement learning methods by more than 4.5 percentage points on the MBPP benchmark and 6.1 percentage points on the Humaneval benchmark. The open-sourced code is publicly available at https://anonymous.4open.science/r/FALCON-3B64/README.md.
Index Terms:
Code generation, Reinforcement Learning, Diverse Feedback.
I Introduction
The development of Large Language Models (LLMs) has significantly advanced automated code generation [1]. Models like CodeLLaMA [2] and DeepSeek-Coder [3], tailored for code-centric tasks, have demonstrated outstanding performance across programming challenges. While LLMs excel in instruction-following through tuning [4], they often misalign with user intent, making feedback-based adjustments critical. For example, InstructGPT [5] leverages reinforcement learning with human feedback (RLHF), and CodeRL [6] uses compilation feedback to refine model performance. Similarly, CompCoder [7] enhances code compilability with compiler feedback, and RLTF [8] offers fine-grained feedback on compiler errors. However, current RL frameworks generate compilation errors and overlook non-differentiable features (e.g. coding style) that affect the performance significantly [4]. To address these challenges, we propose a reinforcement learning system combining long-term and short-term memory feedback. From the global level, long-term memory tracks trends over time for higher-quality code retrieval, while from the local level, short-term memory captures recent errors and immediate feedback. The main contributions of this paper are as follows:
@@ I Introduction
  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to generalize across diverse programming tasks efficiently, adapt to new environments quickly, and handle a wide range of coding challenges with fewer training iterations.
II Related Works
II-A Pre-trained Models for Code Generation
@@
In recent years, pre-trained language models have made significant progress in the field of code generation. Trained on large-scale code corpora, these models have demonstrated powerful code generation capabilities. For example, CodeBERT [9], a model based on an encoder-only architecture, has shown impressive results. With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling [10]. Several models, such as CodeGPT [11], CodeGeeX [12], and DeepSeek-Coder [3], use Causal Language Modelling (CLM) pretraining, while others like CodeT5 [13] and AlphaCode [14] utilize an encoder-decoder architecture. Additionally, models like CodeFusion [15] leverage diffusion-based techniques for code generation. These pre-trained models exhibit great potential in code generation tasks, achieving notable improvements in accuracy through various architectures and training strategies. However, they still face challenges in ensuring the syntactical and functional correctness of the generated code.
II-B Reinforcement Learning on code
@@
Reinforcement learning (RL) is a method that learns optimal strategies through reward signals from interacting with the environment [16]. It excels in sequence generation tasks, such as enhancing performance in translation and summarization models by improving BLEU and ROUGE scores [17]. Unlike traditional natural language processing tasks, code generation is more complex. Beyond syntactical correctness, the generated code must be functionally accurate, meaning it should compile and perform as expected in various scenarios. Passing unit tests is important for verifying correctness, safety, and precision, but not sufficient unless the tests are comprehensive and well-designed. Unit tests must cover diverse use cases and edge cases to ensure the code adheres to standards and meets requirements. For instance, the RL-based code generation fine-tuning framework, CodeRL [6], guides fine-tuning by integrating unit test signals with reinforcement learning. PPOCoder [18] improves on CodeRL by employing Proximal Policy Optimization to refine the approach, while RLTF [8] incorporates specific error information from the code and multi-granularity feedback alongside an online framework for model training. StepCoder [19] enhances the code generation process through compiler feedback and segmental optimization. However, current RL-based methods still face some limitations. Primarily, these methods lack detailed feedback on specific types and distributions of programming errors, which is crucial for identifying patterns, understanding root causes, and creating targeted interventions. This makes addressing recurring issues difficult. Additionally, the corrective mechanisms are not robust enough, often failing to provide timely, specific guidance, hindering effective learning. Additionally, the diversity of input-output examples in existing benchmarks is limited, restricting the model’s ability to adapt and generalize to different or unseen problems [8].
III Problem Setting
We aim to enhance the automated code generation capabilities of Large Language Models (LLMs) by addressing key challenges in accuracy, diversity, error correction, and code quality. Formally, given a high-level specification $D$, the task is to generate a sequence of code tokens $W=\{w_{1},w_{2},\ldots,w_{T}\}$ that maximizes the conditional probability $P(W|D,\theta)$, where $\theta$ represents the model parameters. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\log P(W|D,\theta)\right]$.
@@ III Problem Setting
Challenges. Effective code generation is impeded by several factors. Accuracy requires minimizing syntactical or logical errors in $W$ to ensure correct functionality. Diversity of input-output examples $\mathcal{X}$ is often limited, restricting the model’s ability to generalize across varied programming tasks. Efficient error correction mechanisms are necessary to identify and rectify errors in $W$, ensuring robust performance. Maintaining high code quality, which encompasses adherence to coding standards, style guidelines, and managing code complexity $C(W)$, remains a persistent challenge. Additionally, adaptability refers to the model’s capacity to adapt to new tasks and incorporate feedback for continuous improvement, which is constrained without robust memory mechanisms.
FALCON Framework. To address these challenges, we propose the FALCON framework. The framework leverages both a long-term memory buffer $\mathcal{M}_{\text{long}}=\{(D_{i},W_{i},T_{i},F_{i})\}_{i=1}^{N}$ and a short-term memory buffer $\mathcal{M}_{\text{short}}=\{(D_{j},W_{j},T_{j},F_{j})\}_{j=1}^{M}$ to utilize diverse feedback, enabling fine-tuning of $\theta$. Formally, FALCON seeks to optimize $\theta$ by maximizing a composite reward function $R(W,F)=\alpha T(W)+\beta S(W)+\gamma C(W)+\delta E(W)$, where $T(W)$, $S(W)$, $C(W)$, and $E(W)$ represent unit test, code style, complexity, and error feedback, respectively. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\mathbb{E}_{W\sim P(W|D,\theta)}\left[R(W,F)\right]\right]$.
Assumptions. The effectiveness of FALCON is based on the following assumptions: 1. Exchangeability of the dataset $\mathcal{D}=\{D_{i}\}_{i=1}^{N}$, implying that the order of tasks does not influence model performance; 2. Independence of feedback signals $F$ given the generated code $W$; 3. Sufficient memory capacity in both $\mathcal{M}_{\text{long}}$ and $\mathcal{M}_{\text{short}}$ to store relevant interactions without significant data loss; and 4. Feedback efficacy, ensuring that integrated feedback mechanisms provide meaningful and actionable information to guide the optimization of $\theta$.
Optimization Objective. Given a set of tasks $\mathcal{T}=\{D_{i}\}_{i=1}^{N}$, the goal is to learn optimal parameters $\theta^{*}$ that maximize the expected composite reward across all tasks, formulated as $\theta^{*}=\arg\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}R(W_{i},F_{i})$, where $W_{i}\sim P(W|D_{i},\theta)$ is the generated code for task $D_{i}$, and $F_{i}$ represents the aggregated feedback. The formulation ensures continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
IV Methodology
@@
IV-A Task Definition: Code Generation by LLMs
@@
IV-B MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
We propose the MRLF algorithm 1, which involves random task sampling from the task distribution during training. Previous works have indicated that different data sampling strategies have varying impacts on model information extraction [20]. Consequently, we implement both long and short memory sequences. The long memory strategy stores the solutions generated for each problem and the compiler feedback results, whereas the short memory sequence selects the latest samples and unit test feedback from each current iteration. To address repetitive runtime and compilation errors, the long memory strategy categorizes and stores various errors by their types and records the corresponding error lines, enabling fine-grained reward allocation.
@@ IV-B MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
Task distribution $\mathcal{T}$
Updated model parameters $\theta$
Initialize $LMB$, $SMB$, $\theta$
Populate $LMB$, $SMB$ via few-shot demonstrations
repeat
     Sample batch $\{T_{i}\}$ from $\mathcal{T}$
     for each task $T_{i}$ do
         Code Generation: Generate code $\hat{W}$ and test results; record in $SMB$
         Adaptation: Set $\theta_{i}=\theta$; sample mini-batch from $LMB$, $SMB$
@@
         Update $\theta_{i}$: $\theta_{i}\leftarrow\theta_{i}-\alpha\nabla_{\theta_{i}}L_{\text{inner}}(\theta_{i})$
         Evaluation: Assess $\theta_{i}$ on $T_{i}$; document outcomes
     end for
     Meta-Update: $L_{\text{meta}}=\sum_{i=1}^{N}L_{\text{inner}}(\theta_{i})$; refine $\theta$: $\theta\leftarrow\theta-\beta\nabla_{\theta}L_{\text{meta}}$
until Convergence
return $\theta$
@@
$L_{rl}=-\sum_{t=S_{\text{fine}}}^{E_{\text{fine}}}R_{\text{fine}}(\hat{w}_{t})\log p(\hat{w}_{t}|D,\theta,\hat{w}_{1:t-1})$ (2)
@@
As depicted in Figure 2, we employ a meta-reinforcement learning framework to optimize code generation by integrating both long- and short-term memories for enhanced adaptability. From a global perspective, Long-Term Memory $D_{\text{long}}$ stores historical tasks, generated codes, and feedback to provide valuable context. From a local perspective, Short-Term Memory $D_{\text{short}}$ focuses on recent feedback to enable real-time adjustments. This approach leverages the MAML framework [21] for efficient task adaptation with minimal updates.
@@
IV-C Long-Term Memory Feedback
Given that retrieving information from long-term memory helps improve code quality, we use the FAISS framework [22] to retrieve relevant historical code, feedback, and evaluation scores. Task descriptions and feedback are transformed into embedding vectors and then indexed. During code generation, a query vector from the current task retrieves the top-k most similar historical data to guide the process and avoid past errors. The prompt template is provided in the appendix. Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score. We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\boldsymbol{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS.
During the code generation phase, the current task description $\boldsymbol{t}_{\text{current}}$ and feedback $\boldsymbol{f}$ are transformed into a query vector $\boldsymbol{q}=\phi(\boldsymbol{t}_{\text{current}})$. We compute the similarity between the query vector $\boldsymbol{q}$ and the historical vectors $\boldsymbol{v}_{i}$ using cosine similarity $\text{cos}(\boldsymbol{q},\boldsymbol{v}_{i})$, and retrieve the top-$k$ most similar historical tasks. The retrieval process can be represented as:
$\{(t_{i_{1}},f_{i_{1}},e_{i_{1}}),\dots,(t_{i_{k}},f_{i_{k}},e_{i_{k}})\}=\text{Top-}k\left(\boldsymbol{v}_{i}\mid i=1,2,\dots,n\right)$ (7)